ppb:

The sum of three numbers in order is 57. The equation that represents this is x + (x + 1) + (x + 2) = 57. Which value of x from the set {16, 17, 18, 19} makes the equation true?

ZombieLover:

is this question answered? if so please close thank you :)

Mercury:

to figure out which x value makes the equation true, simply solve for x x + (x + 1) + (x + 2) = 57 the signs are all plus signs which makes distributing easy. you can simply remove the parentheses to get x + x + 1 + x + 2 = 57 three x's together combine to get 3x 1 + 2 make 3 so simplifying, we now have 3x + 3 = 57 subtract 3 from both sides, then divide both sides by 3